JoelTheBoss (joel_the_boss):

Write \(log_6~5\) as a logarithm of base 4.

JoelTheBoss (joel_the_boss):

@sleepyjess Any help? :/

OpenStudy (sleepyjess):

uuuuuhhhhh give me a min

JoelTheBoss (joel_the_boss):

Omg thanks! :D

OpenStudy (anonymous):

A its 100% A

OpenStudy (sleepyjess):

Do you know the change of base formula

JoelTheBoss (joel_the_boss):

@blueSurrency wow you got it right with no options! jk @sleepyjess no I don't :(

OpenStudy (sleepyjess):

darn, that gives an actual numerical value though such as 1.48q4r734rq387...

JoelTheBoss (joel_the_boss):

I know what it isuld I use it here?

JoelTheBoss (joel_the_boss):

how would*

OpenStudy (sleepyjess):

Oh, nevermind, if you don't fully complete it, it just changes the base, ok

OpenStudy (anonymous):

Okay i answered the wrong question so love the sarcasm tho!

OpenStudy (anonymous):

its (log4)5/(log4)6

OpenStudy (sleepyjess):

change of base formula: \(\sf log_bx\\\dfrac{log_dx}{log_db}\)

OpenStudy (sleepyjess):

d is the new base

OpenStudy (sleepyjess):

Please don't give direct answers @blueSurrency

OpenStudy (anonymous):

Thanks for the tip

JoelTheBoss (joel_the_boss):

I see

ganeshie8 (ganeshie8):

you wont like this but i guess below is the simplest method to change base w/o using change of base formula directly : \[\large \log_{\color{Red}{6}}~\color{green}{5}= \log_{\color{red}{4^{\log_46}}}~\color{green}{4^{\log_45}}\] \[\large \color{green}{4^{\log_45}} = \left({\color{red}{4^{\log_46}}}\right)^?\]

OpenStudy (sleepyjess):

O_O I have never seen that before...

OpenStudy (anonymous):

Jess, it's because ganeshie8 gots all those smart brains ;D

OpenStudy (sleepyjess):

he does have all those smart brains...

ganeshie8 (ganeshie8):

<3 idk why but your name reminds me of the most smartest person on openstudy @YanaSidlinskiy

JoelTheBoss (joel_the_boss):

@ganeshie8 O_O How did you do that????

OpenStudy (sleepyjess):

I think we are pretty much all confuzzled about @ganeshie8 's method...

JoelTheBoss (joel_the_boss):

Bayana's_republic

OpenStudy (freckles):

\[\text{ \let } u=\log_6(5) \\ \text{ \to get rid of the base 6 do } 6^u=6^{\log_6(5)} \\ \text{ so we have } 6^u=5 \\ \text{ now we can chose \to solve for u using base 4 instead } \\ \log_4(6^u)=\log_4(5) \\ u \log_4(6)=\log_4(5) \\ \text{ solve for u }\]

OpenStudy (freckles):

that is another way to look at it

OpenStudy (yanasidlinskiy):

LOL, @ganeshie8

JoelTheBoss (joel_the_boss):

ok I have no what you did... but from what I do know is that with what you said \(log_4~(6^u)\) is the same as \(log_4~(5)\) Right?

ganeshie8 (ganeshie8):

that looks neat :)

ganeshie8 (ganeshie8):

sleepyjess's formula is what we use to get the work done @Joel_the_boss : \[\large \sf log_{\color{Red}{b}} \color{green}{x }= \dfrac{log_d \color{green}{x}}{log_d\color{red}{b}}\]

ganeshie8 (ganeshie8):

\[\large \sf log_{\color{Red}{6}} \color{green}{5 }= \dfrac{log_4 \color{green}{5}}{log_4\color{red}{6}}\] thats it! now the right side terms are in base 4 ^

OpenStudy (sleepyjess):

So, if I am reading this right, I used the easy formula and you just showed it differently?

JoelTheBoss (joel_the_boss):

What? That's all? That's easy!!! @ganeshie8 and @sleepyjess You guys saved me hours of struggle...... Wait math cant be this easy... Whats the catch?

ganeshie8 (ganeshie8):

thats right, formlas are meant to be easy and make life simple :)

ganeshie8 (ganeshie8):

*math life ofc

OpenStudy (sleepyjess):

catch? there is no catch... as @ganeshie8 said, formulas are meant to make life simple

OpenStudy (sleepyjess):

I would have never gotten through this module in pre-calc if it weren't for formulas for law of sines and law of cosines...

JoelTheBoss (joel_the_boss):

O_O I love you guys, @ganeshie8 Why arn't you teaching? You should be my teacher!!

OpenStudy (yanasidlinskiy):

*Cough* He's teaching me, not you. <3

OpenStudy (mimi_x3):

hahahah hes got a more prestigious job than teaching And .... Hes one of the the most sought after teachers on OS ;)

OpenStudy (yanasidlinskiy):

Uu....^ He still teaches me. And I know he enjoys it:P

JoelTheBoss (joel_the_boss):

If I was a principal I would pay him twice as much, and give him his own parking spot. xDD Thanks again for the help @ganeshie8 before today a log was a piece of wood. And might wanna check the newest post for questions. :P